### v1

# Real-Time Doppler and Ionospheric Dispersion Correction Techniques for Arbitrary Waveforms Utilizing GPU Compute

Abstract
General requirements for radar digital signal processing are ionospheric distortion and Doppler dispersion correction, which has historically required radar-specific hardware to implement in real time. Although analog solutions are computationally efficient, they often come with system design drawbacks which limit waveform flexibility and can result in an overall increase of system complexity. With improvements in modern general compute systems, real-time digital signal processing is becoming more realizable using non-radar-specific high-performance compute. In this paper, we present an analysis of general Doppler and ionospheric correction algorithms for arbitrary waveforms for radar digital signal processing. We also include considerations for efficient implementation of these algorithms in software, specifically using GPU hardware. This analysis includes metrics of performance such as execution time and error correction accuracy. We also provide recommendations for application in radar signal processing. We identify two algorithms for dispersion correction: an FFT-based method for ionospheric dispersion and a numerical interpolation method via sinc interpolation for Doppler dispersion. Both of these algorithms are able to compensate for dispersion equivalent in accuracy to waveform-specific analytical methods and were able to be performed in real-time on a single NVIDIA H100 GPU. These methods are waveform agnostic and applied directly to the samples, improving system flexibility and making them easy to incorporate into existing software-defined radio systems.
DISTRIBUTION STATEMENT A. Approved for public release. Distribution is unlimited.
This material is based upon work supported by the Department of the Army under Air Force Contract No. FA8702-15-D-0001. Any opinions, findings, conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the Department of the Army.
Daniel J. Vickers is with the Georgia Institute of Technology, Atlanta, GA 30332 USA (e-mail: dvickers6@gatech.edu).
A. H. Mack is with MIT Lincoln Laboratory, Lexington, MA 02421 USA (e-mail: Andrew.Mack@ll.mit.edu).
Idahosa A. Osaretin is with MIT Lincoln Laboratory, Lexington, MA 02421 USA (e-mail: idahosa.osaretin@ll.mit.edu).
I Introduction
Doppler and ionospheric dispersion generate significant distortions for ground-to-space radar and communication systems. Failure to properly correct for these distortions can result in a decrease in the signal-to-noise ratio (SNR) of a received signal, a loss in timing/location accuracy, and more. Doppler time dilation and ionospheric distortion correction can be solved analytically for specific waveforms [1, 2], but general solutions for arbitrary waveforms are computationally expensive and require flexible compute systems. Historically, systems have been built around the use of a limited selection of radar waveforms, like linear frequency-modulated waveforms (LFMs), to limit the solutions required. Enabling flexible waveforms allows systems to optimize waveforms for various properties, such as range resolution, sidelobe control, Doppler tolerance, etc. [3].This drives the desire to build modern radar systems with the ability to process a wide variety of radar waveforms. Dispersive effects must be compensated for, and this drives the desire for modern radar systems to have a general solution for arbitrary waveform correction, which the paper presents a solution for.
The feasibility of compensating for dispersion is complicated by real-time signal processing requirements. For radar agility, it is desirable to process receive signals on a pulse-to-pulse basis, allowing for dynamic and responsive transmit scheduling and processing [4]. Radar signal processing has historically been implemented in radar-specific hardware such as field-programmable gate arrays (FPGAs), analog components, and application-specific integrated circuits (ASICs) due to their high signal processing rates [5]. Although radar-specific hardware provides optimal performance, it often requires longer development times, higher costs, and less flexible system designs. Software flexibility improvements can be obtained using more general computing systems such as central processing units (CPUs) and graphics processing units (GPUs), which could allow for more sophisticated digital signal processing (DSP) solutions [6]. The growth trends of computational power for both CPUs and GPUs have been observed [7]. As on-chip processing power has improved, software-defined radios (SDRs) that utilize general compute systems have become more common in research, communication, astronomy, defense, and hobbyist use [8, 9, 10].
In this paper, we address the issue of real-time radar signal processing on arbitrary waveforms by utilizing general GPU compute to compensate for dispersive effects. We do so by presenting correction methods of ionospheric and Doppler dispersion for SDR. We also explore computational optimizations of these algorithms on a GPU, enabling real-time processing. Documentation of these optimizations and their performance on an NVIDIA H100 GPU are presented here.
We review analytic methods for dispersion compensation of LFMs in sections I-A and I-B. We then present methods of arbitrary ionospheric and Doppler dispersion correction in section II. In order to optimize these algorithms for system flexibility, the methods of arbitrary ionospheric and Doppler dispersion correction presented are numerical methods that can be applied to any discretely-measured signal with no information of the modulation type. We will then compare the arbitrary dispersion correction methods to analytic methods for LFMs in section III, and conclude with a discussion of our study in section IV. A list of all variables used throughout the paper can be found in Table I.
TABLE I: A list of all variables used throughout this paper.
Variable Meaning
$q_{e}$ charge of an electron
$m_{e}$ mass of an electron
$\epsilon_{0}$ permittivity of free space
$c$ speed of light
$f$ frequency
$f_{0}$ waveform start frequency
$E$ total electron content (TEC) along the path of propagation
$K_{2}$ plasma delay constant, linearly proportional to $E$
$B$ waveform bandwidth
$T$ waveform pulse width
NN number of received waveform samples
$v_{r}$ range rate of target
$t$ time
I-A Ionospheric Dispersion Compensation for LFMs
The ionosphere is a relatively plasma-dense layer of the upper atmosphere, starting around an altitude of approximately 50 km. The density of the plasma in this atmospheric layer can fluctuate, but is typically densest in increased sunlight, such as during the day, during the summer, and along the equator[11]. As an electromagnetic signal propagates through the ionosphere, the plasma causes a dispersive effect on the signal. A review of ionospheric effects on radio signals can be found in Appendix A. In brief, ionospheric distortion of a waveform is modeled as a frequency-based time delay of the signal, given as
$\begin{split}\tau(f)=\frac{q_{e}^{2}}{8\pi^{2}m_{e}\epsilon_{0}cf^{2}}E=\frac{K_{2}}{cf^{2}}.\end{split}$ (1)
In (1), $E$ is the total electron content (TEC) along the path of propagation and $K_{2}$ is the plasma delay constant, $q_{e}$ and $m_{e}$ are the charge and mass of an electron, $\epsilon_{0}$ is the permittivity of free space, $c$ is the speed of light, and $f$ is the frequency. The TEC increases when the plasma density of the ionosphere is higher and when the signal propagates through more of the ionosphere. Equation 1 is computed via an approximation that should remain valid for any radar band at HF frequencies or higher. Because the time delay is inversely related to the frequency of the waveform, ionospheric dispersion is particularly impactful on SNR loss and range error at lower-frequency radar bands, such as HF, VHF, UHF, and L. The SNR loss due to dispersion modeled from (1) is also particularly significant for wide-bandwidth signals. Thus, these effects can remain relevant at higher frequency bands given a waveform with sufficient bandwidth.
Ionospheric dispersion correction for waveforms has been performed previously with various methods. High-accuracy methods for correction approximate solutions calculated for specific waveforms [1]. This analytic method can be ideal for cases where known simple waveforms will be transmitted. However, the choice to use an analytic method typically reduces system flexibility. For numerical approximations, another class of methods involves using finite impulse response (FIR) filters with a frequency-dependent response matched to the ionosphere [12, 13, 14]. An overview of modern FIR filtering methods for ionospheric dispersion correction can be found in Ref. [15]. These FIR filters typically operate by representing the ionospheric distortion as a frequency response with three unique input coefficients. All but one of the coefficients are held constant while the variable coefficient is tuned to optimal performance. These methods are computationally efficient, but are not exact. FIR filters are commonly best for applications where one or more parameters of the frequency response are unknown.
To correct for the ionospheric distortion of a waveform, a frequency-dependent correction can be applied by temporally translating the waveform $2\tau(f)$ (to account for the two-way propagation). For continuous waveforms (CWs) with constant frequency, ionospheric effects are reduced to a uniform translation in time (or phase). This problem becomes nontrivial when correcting for waveforms with frequency modulation and more complex frequency characteristics.
A known method to analytically compensate for ionospheric dispersion of LFMs is to apply a correction via predistortion, where the transmitted waveform is inversely distorted in relation to the predicted ionospheric effect. The goal of predistortion is to compensate for the ionosphere on transmit such that the desired waveform is received. This method requires analytic solutions or approximations for any given waveform type. For LFMs, the desired frequency of the transmitted waveform is given as
$f^{3}-\left(f_{0}+\frac{Bt}{T}-\frac{2K_{2}B}{cf_{0}^{2}T}\right)f^{2}-\frac{2K_{2}B}{cT}=0.$ (2)
In (2) $T$, $B$, and $f_{0}$ are the pulse width, bandwidth, and start frequency of the LFM, respectively. One method of solving (2) was presented by Halpin, Urkowitz, and Maron [1]. Their result approximates the frequency as a polynomial expansion of the form
$f(t)=f_{0}+(\mu_{0}+\Delta\mu)t+\gamma t^{2}.$ (3)
There are many ways to determine the coefficients of (3). Here we will note the Pseudo–Chebyshev polynomial expansion method [^1] Note that the results presented in (7) and (8) differ from those presented in Ref. [1] by a factor of 2 on two terms. We concluded that this factor of 2 difference is a minor error and elaborate on this result in Appendix B. from Ref. [1] in which the coefficients $\mu_{0}$, $\Delta\mu$, and $\gamma$ are calculated as
$\mu_{0}=\frac{B}{T},$ (4)
$\Delta\mu=\frac{aB}{2}\frac{T_{1}^{2}+T_{3}^{2}}{T_{1}T_{2}\left(T_{3}-T_{1}\right)}-\mu_{0},$ (5)
and
$\gamma=\frac{aB}{2}\frac{T_{2}}{T_{1}T_{3}\left(T_{3}-T_{1}\right)},$ (6)
where $a=\sqrt{3}/2$ and the coefficients $T_{1}$ and $T_{2}$ are given by
$\begin{split}T_{1}=-\frac{aT}{2}&+\frac{2K_{2}}{c}\bigg{[}\left(f_{0}+\frac{B}{2}\right)^{-2}\\ &-\left(f_{0}+\frac{B(1-a)}{2}\right)^{-2}\bigg{]},\end{split}$ (7)
and
$\begin{split}T_{2}=\frac{2K_{2}}{c}&\bigg{[}2\left(f_{0}+\frac{B}{2}\right)^{-2}-\left(f_{0}+\frac{B(1-a)}{2}\right)^{-2}\\ &-\left(f_{0}+\frac{B(1+a)}{2}\right)^{-2}\bigg{]}.\end{split}$ (8)
We also solved (2) directly via the cubic formula to yield
$f(t)=\frac{1}{3}\left(f_{0}+\frac{Bt}{T}-\frac{2K_{2}B}{cf_{0}^{2}T}-\omega-\frac{\chi}{\omega}\right).$ (9)
Where the coefficients are given by
$\chi=\left(f_{0}+\frac{Bt}{T}-\frac{2K_{2}B}{cf_{0}^{2}T}\right)^{2},$ (10)
$\psi=-2\left(f_{0}+\frac{Bt}{T}-\frac{2K_{2}B}{cf_{0}^{2}T}\right)^{3}-27\frac{2K_{2}B}{cT},$ (11)
and
$\omega=\sqrt[3]{\frac{\psi+\sqrt{\psi^{2}-4\chi^{3}}}{2}}.$ (12)
Note that (2) can have up to three unique solutions. For a practical choice of parameters, (9) is the only purely real solution available. The only physical solutions for frequency are purely real, and therefore the only solution considered is that shown in (9).
Equation (9) cannot be integrated directly to compute the phase of predistorted waveform. However, the phase can be integrated numerically to provide another approximation method for correction of ionospheric effects via predistortion with error related to the method of numerical integration. Both of the predistortion methods shown in (3) and (9) provide useful points of comparison for accuracy of ionospheric correction methods for LFMs.
I-B Background of Doppler Dispersion Correction
Doppler dispersion of a waveform is exactly modeled by the two-way longitudinal Doppler shift where the signal is time dilated such that
$S_{doppler}(t)=S(\alpha t).$ (13)
Here $\alpha=\frac{1+v_{r}/c}{1-v_{r}/c}$, and $v_{r}$ is the range rate of the target [16]. Note that we use the convention where $v_{r}$ is positive when the target is approaching the sensor. For a given signal, $S(t)$, the distorted waveform is given as:
The SNR loss and range error due to (13) is primarily a function of bandwidth and modulation type. As the bandwidth of a waveform increases, so does its Doppler-dispersion error. LFMs are considered relatively Doppler tolerant, but other modulation types will typically incur more error.
For narrow-band signals, it can be sufficient to approximate (13) via frequency conversion (also known as time domain downconversion). In frequency conversion, the signal is multiplied by a tone in the time domain, resulting in an upconversion or downconversion. This algorithm is extremely computationally efficient as it scales with a time complexity [^2] For this paper, we present all computational time complexity and processing error in “Big-O” ($\mathcal{O}$) notation. This notation is common in the fields of physics and computational science, but less so in radar DSP. In this notation, the lead-order complexity of a polynomial describing some performance metric is given. For time complexity, lower order compute times are desirable. For example, a time complexity of $\mathcal{O}(N)$ is considered more optimal than $\mathcal{O}(N^{2})$. of $\mathcal{O}\left(N\right)$. However, the frequency shift is only exact at one frequency, and induces errors for pulses with large bandwidths and/or steep phase slopes [17]. For LFMs, this error scales as $\mathcal{O}\left(B^{2}\right)$, where $B$ is the bandwidth of the signal. For arbitrary waveforms, one typically expects this error to be more impactful.
For any signals with a known analytic representation, (13) can be calculated exactly by resampling the signal according to the target range rate. More specifically, when correcting LFMs, this can be modeled by generating a new LFM with a modified phase slope and pulse width according to (13). This analytic resampling method is typically ideal because of its exact correction and because the time complexity scales as $\mathcal{O}\left(N\right)$. However, when performing DSP, the transmit and receive signals are discretized by sampling the waveform. It can then become desirable to pass these samples between software interfaces in order to simplify the overall software complexity of the system. In these cases where no analytic representation of the signal is available, one is forced to use the general solution to approximate the resampling numerically. There are many known ways to compute this numerical resampling. For this paper, we consider the following methods:
  • Polynomial Interpolation: A linear, cubic, higher order, or spline interpolation function is used. While this method is computationally efficient, typically scaling as $\mathcal{O}\left(N\right)$, it results in a loss of power in the interpolated signal due to interpolation error. Higher sample rate signals reduce this error, but raise the computational load.
  • FFT P/Q Resampling: The Fourier transform of the signal has terms removed or added followed by an inverse Fourier transform is taken. This results in the exact resampling assuming a Nyquist-limited waveform. This method can be computationally expensive due to requiring discrete Fourier transforms (DFTs) of non-ideal numbers of samples. This leads to variable sample rates that depend upon the specific number of sample used, and scales as $\mathcal{O}\left(N\log{N}\right)$ in the ideal case. This method is also only exact to an integer number of samples added or removed. A brief exploration of the accuracy and performance of FFT resampling can be found in Ref. [18].
  • Sinc Interpolation: Allows exact resampling of Nyquist limited signals. The exact form of sinc interpolation grows as $\mathcal{O}\left(N^{2}\right)$ in time, but windowing the filter to a size of $N_{window}$ samples results in a time complexity of $\mathcal{O}\left(N\times N_{window}\right)$. The choice of features such as taper window size, taper, and initial sampling rate determine the performance of this algorithm. An overview of Whittaker–Shannon interpolation and general interpolation of Nyquist-limited signals can be found in Ref. [19].
II Methods
We sought a general solution to previous dispersion correction methods by leveraging modern hardware to enable algorithms that were previously too computationally intense. In this section, we present arbitrary methods of Doppler and ionospheric dispersion correction. In Section II-A, we present a numeric method for correcting ionospheric dispersion via a strict time delay from (1) which is applied in the frequency domain. In Section II-B we use (13) to represent Doppler dispersion as an interpolation and explore various methods of interpolation valid for GPU compute.
II-A Ionospheric Dispersion Compensation Via FFT Phase Correction
In this section, we propose a method for numerically determining the ionospheric distortion correction of a transmitted or received waveform in real time directly from discrete samples. Recall from the time-shifting property of the Fourier transform that any time shift in the time domain is represented as a phase shift in the frequency domain. One can apply a frequency-dependent phase shift for each frequency bin using the temporal translation from (1). Thus, using a time delay of $2\tau(f)$ (to account for two-way ionospheric dispersion) one can conclude that the ionospheric effect on a transmitted waveform, $S_{Tx}$, due to two-way propagation will result in a received waveform, $S_{Rx}$, of
$\begin{split}S_{Rx}&=\mathcal{F}^{-1}\left[\mathcal{F}(S_{Tx})\exp\left(2\pi i(2\tau)f\right)\right]\\ &=\mathcal{F}^{-1}\left[\mathcal{F}(S_{Tx})\exp\left(\frac{4\pi iK_{2}}{cf}\right)\right].\end{split}$ (14)
On its own, (14) is a convenient algorithm to directly apply ionospheric distortion to arbitrary waveforms in software. Equation (14) can be applied to the transmitted signal to use as the matched filter during pulse compression. Solving for the inverse of (14) results in a method to remove the ionospheric distortion accrued during propagation and is identical to applying a temporal shift of $2\tau(f)$. The inversion of (14) yields
$\begin{split}S_{Tx}=\mathcal{F}^{-1}\left[\mathcal{F}(S_{Rx})\exp\left(\frac{-4\pi iK_{2}}{cf}\right)\right].\end{split}$ (15)
A similar method to (15) was applied by Peters, Schroeder, and Romero-Wolf [20] for post-processing of satellite data when determining the TEC values of Europa’s atmosphere. Here (15) is intended to correct for the ionosphere in real time given a known TEC value. The error of this algorithm is linearly related to the frequency resolution in the frequency domain. By the Nyquist-Shannon sampling theorem, this means that the error scales as $\mathcal{O}(t_{range})$, where $t_{range}$ is the total length of time of the range window being processed.
In practice, the Fourier transformations used in (15) must be done discretely as Fast Fourier Transforms (FFTs) and inverse FFTs (iFFTs). In the past, it has been too slow to perform the FFT and iFFT in real-time applications on high-bandwidth signals. Modern improvements in hardware and algorithms have made this computation achievable to implement digitally in real time.
Compared to using a polynomial expansion technique, this proposed method offers several benefits. The technique shown in (15) is a direct calculation of the frequency dependent phase shift. This results in a few lines of code, allowing a first step in prototyping with little chance for errors. This technique can distort a waveform on transmit (predistortion) or on receive (post-distortion), and is valid for arbitrary waveforms. The algorithm’s only requirement is that the waveform is sampled within the Nyquist limit. When an analytical solution is inaccessible, (15) can be used as a general numerical method.
The primary drawback of (15) is that it is more computationally expensive in cases where an analytic method can be used. This high computational cost is primarily due to the use of the FFT and iFFT algorithms. The time complexity of (15) scales as $\mathcal{O}\left(N\log{N}\right)$ when the number of samples is exactly a power of two ($N=2^{n}\text{ for }n\in\mathbb{Z}$) and scales as much as $\mathcal{O}\left(N^{2}\right)$ in other cases. The only way to mitigate this cost is with more processing power. The increased processing requirement is why (15) is an excellent candidate for parallelization on a GPU. The computational work done in the frequency domain can also be bundled with other calculations that require FFTs in the signal processing chain and benefit from the efficiency of the shared use of FFT compute time. An example flow diagram of how one could combine pulse compression with (15) for post-distortion is shown in Fig. 1. These optimizations can make our method viable in SDRs to be used for real-time processing in radars and communication systems.
Figure 1: Overview of process of performing pulse compression on the GPU with ionospheric distortion correction.
II-B Numeric Doppler Dispersion Correction Via Interpolation
In this section, we present an overview analysis of various Doppler-dispersion correction methods, all of which seek to approximate the effect of (13) by numerically interpolating the transmitted signal. Table II lists the features of the Doppler dispersion correction methods considered in this work [^3] We determined the error drop off for windowed Whittaker–Shannon interpolation as well as frequency conversion numerically. The data presented in the table reflects the drop off for LFMs.. Of note, we highlight the Whittaker–Shannon interpolation (also known as sinc interpolation) and FFT P/Q resampling methods, both of which are exact for Nyquist-limited signals.
TABLE II: Comparison of features of various Doppler dispersion correction/discrete resampling methods.
Method Name Time Complexity Accuracy Arbitrary Waveforms
Whittaker–Shannon Interpolation $\mathcal{O}\left(N^{2}\right)$ Exact Yes
Windowed Whittaker–Shannon Interpolation $\mathcal{O}\left(N\times N_{window}\right)$ $\mathcal{O}\left(N_{window}^{-2}\right)$ Yes
Frequency Conversion $\mathcal{O}\left(N\right)$ $\mathcal{O}\left(B^{2}\right)$ (for LFMs) Yes
FFT P/Q Resampling Variable Exact for Even Whole Indices Yes
FFT P/Q Resampling w/ Frequency Conversion Variable Exact for Even Whole Indices Yes
Analytical Resampling $\mathcal{O}\left(N\right)$ Exact No
Linear Interpolation $\mathcal{O}\left(N\right)$ $\mathcal{O}\left(N^{-1}\right)$ Yes
Whittaker–Shannon interpolation is given by (16):
$x(t)=\sum_{n=-\infty}^{\infty}x[n]\operatorname{sinc}\left(\frac{t-n\Delta t}{\Delta t}\right).$ (16)
In (16), $\Delta t$ is the sample period of the discrete signal. (16) is general and does not require upsampling to achieve exact performance. Unfortunately, the time complexity of Whittaker–Shannon interpolation scales as $\mathcal{O}\left(N^{2}\right)$. For more optimal interpolation rates, Whittaker–Shannon requires a window where only $N_{window}$ neighboring points are used in the summation. This was investigated in Ref. [21].
FFT P/Q resampling is mathematically identical to (16), but is performed in the frequency domain. In literature, FFT P/Q resampling can also be refered to as Whittaker–Shannon interpolation—we define these as separate algorithms for clarity. Via the Fourier transform, the sinc filter used in Whittaker–Shannon interpolation becomes a box filter, zeroing out or adding zeroed edge samples near $f=\pm\frac{f_{s}}{2}$. The size of this box filter is $\frac{N}{\alpha}$. After appending or removing the appropriate number of samples, an iFFT is performed using the same number of samples as the box filter, which interpolates the signal. This interpolation algorithm is common in image processing [22] as it is relatively simple to implement and can utilize optimized FFT libraries.
A key drawback of FFT P/Q resampling vs. Whittaker–Shannon interpolation is that FFT P/Q resampling is only exact when $N-\frac{N}{\alpha}$ is an even integer. FFT P/Q resampling can be combined with frequency conversion for fine-tune adjustment between even-integer interpolant results. This is done by first performing FFT P/Q resampling to the nearest velocity that removes an even number of samples, and then performing frequency conversion to shift the remaining frequency difference. This presents a trade space for the optimal time dilation method that depends on the center frequency, bandwidth, signal length, and modulation.
III Analysis
In this section, we present analysis of the signal processing accuracy and rate of each algorithm. To compare the accuracy of each dispersion correction algorithm, we compared the arbitrary signal processing methods presented in section II with those presented in sections I-A and I-B. In order to compare algorithms directly, we performed an analysis of the signal processing accuracy via matched filtering of LFMs. To demonstrate the computational rates of the methods considered, all algorithms in this paper were implemented in CUDA 12.3 and performed on the NVIDIA H100 GPU with a $3.25~\text{GHz}$ CPU processor connected via PCIe 5.0. All calculations performed in both forms of analysis used a single-precision complex float, otherwise known as a “complex64”, for each I/Q sample. The signal processing accuracy and timing performance for the ionospheric dispersion correction are presented in section III-A. A similar analysis of the Doppler dispersion algorithms is presented in section III-B.
III-A Ionospheric Correction Performance Analysis
We tested (15) against the Pseudo–Chebyshev approximation method shown in (3). Each method was compared against truth data calculated via the cubic solution from (9) using trapezoidal integration to determine the phase. Results from one of our sample test cases can be seen in Fig. 2. As can be seen in this example, the FFT approximation method and the Pseudo–Chebyshev method resulted in losses on the order of $0.01~\text{dB}$ SNR. The difference between each method resulted in peaks separated by $0.001~\text{dB}$. This analysis indicated that (15) is nearly identical in accuracy to the Pseudo–Chebyshev polynomial approximation method for the parameters used to generate Fig. 2.
Figure 2: Comparison of the FFT-based ionospheric compensation method with both the undistorted waveform (labeled LFM), the Pseudo–Chebyshev polynomial approximation (labeled POLY), our FFT based method (labeled FFT), and our exact frequency from the cubic formula shown in Eq. 9 integrated via trapezoidal integration (labeled CUBIC). This test was run with $f_{0}=413~\text{MHz}$, $B=18~\text{MHz}$, a sample rate of $f_{s}=2.048~\text{GHz}$, $E=100e16~\text{electrons}/\text{m}^{2}$, and $T=100~\mu\text{s}$. In this particular example, $\text{CUBIC}\circledast\text{FFT}$ curve compares the FFT-based correction method with truth and lost $0.01~\text{dB}$ SNR. For this example, the Pseudo–Chebyshev polynomial expansion method compared similarly as close as $0.001~\text{dB}$, indicating that the FFT method is on par with the accuracy of the Pseudo–Chebyshev polynomial method.
The choice of GPU hardware allowed us to parallelize the FFT, iFFT, and the phase shifting of each sample in the receive window. Figure 3 shows the timing performance result of running (15) on a realistic waveform. Notably, Fig. 3b shows timing performance excluding the FFT and iFFT compute times. We separate these test cases to demonstrate the efficiency that can be gained from bundling the ionospheric correction with other digital signal processing methods that require computations in the frequency domain.
Figure 3: Histograms showing the results from $10^{7}$ time trials where we performed the algorithm shown in (15). This test was run with $f_{0}=413~\text{MHz}$, $B=18~\text{MHz}$, a sample rate of $f_{s}=2.048~\text{GHz}$, $E=100e16~\text{electrons}/\text{m}^{2}$, $T=100~\mu\text{s}$, and had $2^{19}$ total I/Q samples. Figure 3a shows the rates when performing ionospheric correction during pulse compression, which averaged about $46~\mu\text{s}$ to complete and correlates to a rate of approximately $11.3~\text{Gigasamples}/\text{s}$. Figure 3b excludes the FFT and iFFT times and averages approximately $2200~\text{ns}$, which correlates to about $240~\text{Gigasamples}/\text{s}$.
As can be seen in Fig. 3, on our hardware we were able to complete the ionospheric correction considered in about $46~\mu\text{s}$; nearly $44~\mu\text{s}$ of which are from performing the required FFTs. The signal processing time to compensate for ionospheric dispersion via (15) is significantly lower than the pulse width of the waveform itself, even at a relatively high sampling rate of $f_{s}=2.048~\text{GHz}$. This efficiency demonstrates the viability of (15) for use in real-time DSP.
III-B Performance Analysis of Doppler Dispersion Correction
An overview of the accuracy of each method from Table II is summarized in Fig. 4. For low velocities, the frequency conversion method has the lowest computational load and performs well. Even when at even sample removal, the FFT P/Q resampling with or without the tone has nearly equivalent accuracy to Whittaker–Shannon interpolation at the cost of variable compute times. At non-integer interpolants, the sinc/Whittaker–Shannon method performs the dilation with the least error.
Figure 4: Comparison of error as a function of target velocity for the various Doppler dispersion correction methods considered in Table II. The waveform used here had $f_{0}=411~\text{MHz}$, $B=18~\text{MHz}$, and $T=500~\mu\text{s}$. The waveform was originally sampled at a rate of $f_{s0}=2~\text{GHz}$, and was downconverted to $f_{s}=200~\text{MHz}$. Notice that the the error for FFT P/Q resampling is harmonic, with no error at velocities that require an even integer number of samples be padded and maximal error for velocities that require an odd integer number of samples be added. Note that for the sinc interpolation loss shown here, a window size of $N_{window}=25~\text{samples}$ was used as well as a lookup table (LUT) that was sampled such that for $\operatorname{sinc}(x)$, $\Delta x=1/100$ ($100~\text{values}/\text{sample}$). Refer to Appendix C for details on the implementation of LUTs.
For FFT P/Q resampling, the signal can be resampled during an FFT by padding the signal with an even number of samples. When the number of interpolants, $N_{I}$, is a power of two, FFT P/Q resampling completes in $\mathcal{O}\left(N_{I}\log{N_{I}}\right)$ time, but the time complexity is much more sophisticated and almost always less optimal in other cases. The variable time complexity is demonstrated in Fig. 5 where three separate clusters of Doppler dispersion correction rates can be observed, each cluster correlated to a range of target velocities. This means that FFT P/Q resampling has a pseudo-velocity-dependent time complexity, which can result in unpredictable signal processing rates. The test case used in Fig. 5 is considered relatively stressful in terms of typical real-time DSP, and the required compute could be considered viable for some systems. While the simplicity of implementing FFT P/Q resampling is desirable, the variable time complexity is likely only be valid for systems with low sample rates or range extents.
Figure 5: Doppler dispersion correction rates for $10^{6}$ trials of FFT P/Q Resampling with a waveform such that $f_{0}=420~\text{MHz}$, $B=18~\text{MHz}$, $T=500~\mu\text{s}$, and $N=2^{19}$. To demonstrate the variable time complexity, we also varied the target velocity by using a uniform distribution between $v=0\text{ to }5~\text{km}/\text{s}$. Notice three separate clusters of completion times, each related to a different range of velocities. The left-most cluster is a sharp peak with times $\leq 10~\mu\text{s}$ and corresponds to about $v<1500~\text{km}/\text{s}$ where no samples were added (no work was done). The right-most peak was for velocities such that about $1500<v<4500~\text{km}/\text{s}$, where only 2 samples were removed. The final middle cluster was the remaining trials during which 4 samples were removed. The average completion rates for the left, center, and right cluster are about $6.5~\text{Terasamples}/\text{s}$, $670~\text{Megasamples}/\text{s}$, and $260~\text{Megasamples}/\text{s}$ respectively.
Considerations for efficient implementation of Whittaker–Shannon interpolation are presented in Appendix C. Our optimal implementation of the algorithm is presented in Alg. 2 and resulted in interpolation rates shown in Fig. 6 of about $11.7~\text{Gigasamples}/\text{s}$. For a radar transmitting at 10% duty factor, this would mean a rate of $117~\text{Gigasamples}/\text{s}$ could be sampled on receive.
Figure 6: timing performance measurements of Whittaker–Shannon interpolation times on an NVIDIA H100 GPU for $10^{7}$ time trials. The input waveform contained $N=2^{19}~\text{samples}$ total with a window size of $N_{window}=25~\text{samples}$. Interpolation was achieved at an average of $45~\mu s$, or a rate of about $11.7~\text{Gigasamples}/s$.
IV Conclusion
In this paper, we investigated potential algorithms for ionospheric and Doppler dispersion correction of arbitrary waveforms. The goal of this analysis was to identify algorithms which are viable for real-time DSP on GPU hardware, allowing for simpler system designs. We also desired flexible algorithms by implementing numerical approximation methods that could be applied directly to discrete signals with no requisite knowledge of the transmitted waveform.
For ionospheric dispersion correction, we proposed an FFT-based method which was presented in (15). This algorithm is computed directly from the ionospheric delay shown in (1). The primary benefits of this method are that it requires no previous information of the input waveform to compute, is valid for both predistortion and post-distortion, is valid for arbitrary waveforms, and is straightforward to implement. The major drawback of this algorithm is the relatively high computational cost, which can only be mitigated with more powerful parallel compute hardware. This algorithm also has an SNR loss of $\mathcal{O}(t_{total})$, which means that the accuracy is reduced for smaller range window sizes. The signal processing accuracy of (15) was shown in Fig. 2 on LFMs. This demonstrated that the algorithm was approximately as accurate as the Pseudo–Chebyshev polynomial expansion method, which indicates a suitable accuracy for radar DSP. In Figs. 3a and 3b we provided timing performance of an implementation of (15) on an NVIDIA H100 GPU. These timings demonstrate that this method can be viable for application in real-time DSP given appropriate selection of hardware and parallelization.
We also presented a comparison of various Doppler dispersion correction methods via time dilation. We highlighted the methods of FFT P/Q resampling and windowed Whittaker–Shannon interpolation, given in (16), as potentially viable candidates for Doppler dispersion correction in SDRs. Figure 4 presents the accuracy of various Doppler compensation algorithms considered. The data processing rates of FFT P/Q resampling and sinc interpolation on GPU hardware are shown in Figures 5 and 6 respectively. Both algorithms were demonstrated to have sufficient accuracy to be considered for use in radar DSP. The signal processing rates of windowed Whittaker–Shannon interpolation were much more consistent and rapid than those of FFT P/Q resampling. While windowed Whittaker–Shannon interpolation was shown to be valid for radar DSP, the variable processing rates of FFT P/Q resampling mean that it is likely only viable for systems with low sampling rates or range window sizes until modern GPU hardware improves. The major benefit of FFT P/Q resampling is its simple implementation and optimization, which results in relatively little code. On the other hand, optimization of efficient windowed Whittaker–Shannon interpolation is more challenging. To reduce this challenge, we provide considerations for optimizing windowed Whittaker–Shannon interpolation in Appendix C. Despite these considerations, the ability to implement windowed Whittaker–Shannon interpolation on a GPU simplifies system design time relative to radar-specific hardware. We conclude that windowed Whittaker–Shannon interpolation is the most-viable candidate for real-time radar DSP considered in this work and would be sufficient for many modern real-time DSP applications.
Appendix A Ionospheric Effects on Radio Signals
Ionospheric effects can be modeled as temporal shifts of individual frequencies of a signal. According to Davies [23] the ionospheric index of refraction, $n_{G}$, is represented as shown in (17):
$n_{G}=1+\frac{q_{e}^{2}}{8\pi^{2}m_{e}\epsilon_{0}f^{2}}N_{e}.$ (17)
Here, $N_{e}$ is the ionospheric electron density and $\epsilon_{0}$ is the permittivity of free space. Thus, the group velocity of an ionospheric radio wave, $V_{G}$, is given as (18):
$\begin{split}V_{G}=\frac{c}{n_{G}}&=c\left(1+\frac{q_{e}^{2}}{8\pi^{2}m_{e}\epsilon_{0}f^{2}}N_{e}\right)^{-1}\\ &=c\left(1-\frac{q_{e}^{2}}{8\pi^{2}m_{e}\epsilon_{0}f^{2}}N_{e}\right)+\mathcal{O}\left(f^{-4}\right).\end{split}$ (18)
We can truncate the Taylor series expansion in (18) at $\mathcal{O}\left(f^{-2}\right)$ for sufficiently high values of $f$. To provide context, in SI units this truncation approximation is typically valid for all waveforms such that $f>>6~\text{MHz}$ in Earth’s atmosphere. We can use the group velocity to integrate in time and determine the one-way separation $R_{e}$ between the radar and the target as
$R_{e}=\int_{\Delta t}V_{G}dt=c\Delta t-\frac{q_{e}^{2}}{8\pi^{2}m_{e}\epsilon_{0}f^{2}}\int_{\Delta t}cN_{e}dt.$ (19)
Making the substitution $ds=cdt$ to integrate along the propagation path of the pulse yields
$R_{e}=c\Delta t-\frac{q_{e}^{2}}{8\pi^{2}m_{e}\epsilon_{0}f^{2}}\int_{s}N_{e}ds.$ (20)
The propagation time given no ionospheric distortion effect, $\Delta t$, would produce a range of $R_{0}=c\Delta t$. We solve for the one-way temporal shift $\tau$ of an ionosphere distorted waveform to be
$\begin{split}\tau(f)&=\frac{R_{0}-R_{e}}{c}=\frac{q_{e}^{2}}{8\pi^{2}m_{e}\epsilon_{0}cf^{2}}\int_{s}N_{e}ds\\ &=\frac{q_{e}^{2}}{8\pi^{2}m_{e}\epsilon_{0}cf^{2}}E=\frac{K_{2}}{cf^{2}}.\end{split}$ (21)
Appendix B Correction to Pseudo–Chebyshev Polynomial Expansion Method
The original Pseudo–Chebyshev method presented in Ref. [1] slightly differed from the coefficients we presented in (7) and (8). The original forms of $T_{1}$ and $T_{2}$ were given as:
$\begin{split}T_{1}=-\frac{aT}{2}&+\frac{4K_{2}}{c}\bigg{[}\left(f_{0}+\frac{B}{2}\right)^{-2}\\ &-\left(f_{0}+\frac{B(1-a)}{2}\right)^{-2}\bigg{]},\end{split}$ (22)
and
$\begin{split}T_{2}=\frac{4K_{2}}{c}&\bigg{[}2\left(f_{0}+\frac{B}{2}\right)^{-2}-\left(f_{0}+\frac{B(1-a)}{2}\right)^{-2}\\ &-\left(f_{0}+\frac{B(1+a)}{2}\right)^{-2}\bigg{]}.\end{split}$ (23)
The values of $T_{1}$ and $T_{2}$ shown in (22) and (23) only differ by using a factor of $4$ instead of $2$ with each $K_{2}$ term. We originally noticed this difference numerically by comparing parameters for the FFT ionospheric correction method presented in (15). We then repeated the original derivation from Ref. [1] as well as analytically derived an exact solution to (2), which we show in (9). We note that in Ref. [1] a substitution of a factor of $2$ is made for a factor of $4$, about which the authors state was determined “by simulation”. Without more information, we are unable to repeat their simulations to determine the validity of that substitution. However, when we rederived the Pseudo–Chebyshev method polynomial expansion, we arrived at the solutions shown in (7) and (8), which the FFT ionospheric correction method agrees with.
Figure 7: Frequency residuals of the Pseudo–Chebyshev method with our exact solution shown in (9) for a given LFM with $B=20~\text{MHz}$, $f_{c}=400~\text{MHz}$, and $T=0.1~\text{ms}$. Here, the corrected residual is the difference returned when using the coefficients in (7) and (8); the uncorrected curve using the coefficients in (22) and (23). Note that in this case the corrected Pseudo–Chebyshev method resulted in the lower error by over a full order of magnitude.
To further verify this change, we compared the original formulation of the Pseudo–Chebyshev method with our corrected coefficients, which is shown in Fig. 7. As shown in the plot, our corrected coefficient values applied for the given frequency resulted in a difference over a full order of magnitude lower than with the uncorrected method. This difference resulted in about $1~\text{dB}$ of SNR loss when cross correlated with our cubic solution. Based upon the errors from our exact analytic solution, comparison with our FFT-based ionospheric correction method, and analytic derivations, we conclude that the original factor of 2 difference presented in Ref. [1] was a slight error, which we have corrected here.
Appendix C Algorithmic Optimizations for Sinc Interpolation on a GPU
For optimal performance, we elected to implement windowed Whittaker–Shannon interpolation on an NVIDIA H100 GPU connected to a $3.25~\text{GHz}$ CPU via PCIe 5.0. Our implementation was written in CUDA C++. Implementation on a GPU allowed us to parallelize a large portion of the computations. Here we will discuss three potential optimizations that were considered, their effect on performance in the H100, and considerations for use on different hardware.
Figure 8: Timing performance for this paper’s implementation of windowed Whittaker–Shannon interpolation times on an NVIDIA H100 GPU utilizing 32 blocks per kernel. Each plot demonstrates kernel completion times for four different implementations considered and ran for $10^{7}$ time trials. The input waveform contained $N=2^{19}~\text{samples}$ with a window size of $N_{window}=25~\text{samples}$. Figure 8a shows the interpolation time with no optimizations made outside of basic parallelization, which averaged completion around $150.5~\mu\text{s}$. Figure 8b displays the timing data if we first pre-load the input array into shared memory (tiling), which averaged completion around $159~\mu\text{s}$. Figures 8c and 8d present the timing performance when a LUT was included. The LUTs were size 3000 and 300 for the Figures 8c and 8d respectively. These kernels averaged completion in about $60.5~\mu s$ for the LUT size of 3000 and $59.3~\mu\text{s}$ for the LUT size of 300.
The primary optimization made during the implementation of this code was parallelization on a GPU. The standard method of parallelizing code on a GPU is to assign each executing thread to an index in the output array. This procedure is not only an efficient way to organize the workload on any given thread, but also assigning each thread a unique output memory address guarantees thread safety during the kernel call. In the case of windowed Whittaker–Shannon interpolation, each thread is responsible for the single summation of a window and writing the resultant interpolant to the output array. Recall from Table II that for an output array of $N$ elements, and a window size of $N_{window}$ elements, windowed sinc interpolation requires $\mathcal{O}\left(N\times N_{window}\right)$ time to complete. If the number of threads in the kernel meets or exceeds $N$, then the parallelized computation on a GPU reduces to $\mathcal{O}\left(N_{window}\right)$ time. For clarity, pseudo-code of this implementation can be seen in Algorithm 1 and we visually represented this parallelization in Fig. 9.
Algorithm 1 Parallelized Windowed Whittaker–Shannon Interpolation
1:procedure Sinc Interpolation Kernel
2:  $\text{index}=\text{index of thread in grid}$
3:  $\text{stride}=\text{total threads in grid}$
4:  for $\text{index}<\text{number of samples}$ do
5:   $\text{center}=\text{nearest uninterpolated point to index}$
6:   $\text{sum}=0$
7:   for input samples in window around center do
8:     sinc value=calculated $\text{sinc value}$$\text${calculated sinc value}
9:     $\text{sum}=\text{sum}+\text{input sample}\times\text{sinc value}$
10:   $\text{output array}[\text{index}]=\text{sum}$
11:   $\text{index}=\text{index}+\text{stride}$
Figure 9: A visual representation of the data accesses during a summation of our parallelized implementation of windowed Whittaker–Shannon interpolation. In this parallelization, every thread in the GPU kernel is allocated to an output index $y_{n}$. For a given thread such that $n=d$, the output index will be $y_{d}$. That worker thread identifies the nearest samples to it in the original waveform, shown as $x_{c}$, which will be the center of the window. The thread then considered all values of $x_{n}$ in the window, which range from $x_{c-(N_{window}-1)/2}$ to $x_{c+(N_{window}-1)/2}$. After the summation, the total is written to the output index.
After parallelization on the GPU, the highest-yielding optimization for our implementation of windowed sinc interpolation was to reduce the time cost to compute the values of the $\operatorname{sinc}$ function. The sine function (and therefore the $\operatorname{sinc}$ function) is generally approximated via a Taylor series expansion, which is slow to compute. One can save compute time by precalculating the values of the sinc function for use in the summation and storing them in an array known as a lookup table (LUT). For fast memory access that is visible to all threads in a block, we store the LUT in shared memory on a GPU. The accuracy of the sinc values from the LUT will depend upon the rate at which the $\operatorname{sinc}$ function was sampled, with higher sample rates leading to more accuracy and larger LUTs. The compute time required to initialize this LUT into shared memory depends on the LUT size. Note that since $\operatorname{sinc}$ is an even function, only the positive values of $\operatorname{sinc}$ are required, which saves LUT population time and shared memory. On a GPU, creating a LUT in shared memory offers a significant performance increase during the summation and only comes at the cost of minor wait times to synchronize the threads after populating the LUT. In our case, addition of a LUT improved performance by as much as $2.5\times$ faster interpolation rates when using a LUT size that provided near-exact accuracy. We present pseudo-code for this optimization in Algorithm 2.
Algorithm 2 Parallelized Windowed Whittaker–Shannon Interpolation w/ Lookup Table
1:procedure Sinc Interpolation Kernel
2:  $\text{index}=\text{index of thread in grid}$
3:  $\text{stride}=\text{total threads in grid}$
4:  $\text{sinc LUT}=\text{allocated shared memory}$
5:  for element in sinc LUT do
6:   $\text{sinc LUT}[\text{element}]=\text{precalculated sinc value}$
7:  synchronize threads
8:  for $\text{index}<\text{number of samples}$ do
9:   $\text{center}=\text{nearest uninterpolated point to index}$
10:   $\text{sum}=0$
11:   for input samples in window around center do
12:     $\text{sinc value}=\text{value in sinc LUT}$
13:     $\text{sum}=\text{sum}+\text{input sample}\times\text{sinc value}$
14:   $\text{output array}[\text{index}]=\text{sum}$
15:   $\text{index}=\text{index}+\text{stride}$
The final optimization we consider here is reduction in the number of read accesses to global memory from the input array of samples. Each thread in the kernel is responsible for $N_{window}$ global memory accesses—one for each element in the summation. Fortunately all threads within a block read from a highly-overlapping region of the input array. Therefore, global memory access can be reduced by allocating a region of shared memory for each block in the kernel, and then having each thread perform a single global memory read to initialize the input array into shared memory. This is a method of preloading the global memory into the block’s shared memory, referred to as “tiling”. Typically tiling is utilized to accelerate memory accesses in a 2D array (with each 2D region being a tile), but it can be applied in our 1D case. The major drawback of the use of tiling is the time cost of thread synchronization that must occur each time the shared memory is updated with a new region of input samples. We found that for our typical choice of interpolation parameters, tiling resulted in about a $9-10~\mu\text{s}$ slow-down in computational time. Because the number of memory accesses to the input array scales as $\mathcal{O}\left(N_{window}\right)$, tiling should become optimal for large enough choice of $N_{window}$, and may be optimal on another choice of hardware where global memory accesses are slower. Because tiling was slightly slower for our use case and for simplicity, we chose to exclude this potential optimization from our implementation of Whittaker–Shannon interpolation. Algorithm 3 presents pseudo-code for our version of sinc interpolation with tiling.
Algorithm 3 Parallelized Windowed Whittaker–Shannon Interpolation w/ Tiling
1:procedure Sinc Interpolation Kernel
2:  $\text{index}=\text{index of thread in grid}$
3:  $\text{stride}=\text{total threads in grid}$
4:  $\text{preloaded values}=\text{allocated shared memory}$
5:  for $\text{index}<\text{number of samples}$ do
6:   for element in preloaded values do
7:     $\text{preloaded values}[\text{element}]=\text{global value}$
8:   synchronize threads
9:   $\text{center}=\text{nearest uninterpolated point to index}$
10:   $\text{sum}=0$
11:   for input samples in window around center do
12:     $\text{input value}=\text{value from preloaded values}$
13:     sinc value=calculated $\text{sinc value}$$\text${calculated sinc value}
14:     $\text{sum}=\text{sum}+\text{input sample}\times\text{sinc value}$
15:   $\text{output array}[\text{index}]=\text{sum}$
16:   $\text{index}=\text{index}+\text{stride}$
17:   synchronize threads
Note that the implementations of a LUT and tiling both require shared memory allocation that is related to $N_{window}$ in size. It is possible that tiling and the use of a LUT will be incompatible on devices with sufficiently limited shared memory and/or a sufficiently large choice of $N_{window}$. Based on our timing performance and because the number of sinc value calculations scales as $\mathcal{O}\left(N_{window}\right)$, it seems likely that the implementation of a LUT will most-likely be optimal in any situation where a LUT and tiling are mutually exclusive.
Most further optimizations for a GPU implementation of sinc interpolation are much more specific to the given hardware used, such as maximal occupancy requirements. These optimizations are more sophisticated than we intend to include here. A comparison of the efficiency of each considered optimization on our hardware limited to 32 blocks per kernel is shown in Fig. 8.
For completeness, we note that a great deal of work has been done on efficient Whittaker–Shannon interpolation by padding elements in the frequency domain via the use of DFTs [24, 25]. For efficient processing, these methods require that the number of interpolants be proportional to the number of input samples by a power of 2. This condition is compatible with applications where the number of interpolants, $N_{I}$, is selectable such as image processing. When this condition is met, Whittaker–Shannon interpolation is achievable with time complexities of $\mathcal{O}\left(N_{I}\log{N_{I}}\right)$. Interpolation for Doppler dispersion correction requires a specific choice of $N_{I}$ that depends upon $N$ and the target velocity. The variable number of interpolants results in variable time complexity to complete the required FFTs. For our use case, these methods resulted in an unpredictable signal processing rate that was consistently slower than all implementations shown in Fig. 8—in some cases by a factor of $10\times$ slower rates. These trade offs are extremely similar to considerations for the FFT P/Q resampling method. We conclude that for radar signal processing, performing the interpolation in the time domain directly from (16) is often more efficient and always more predictable in terms of kernel completion times.